OpenStudy (anonymous):

Evaluate: 32 + (6 − 2) • 4 − 6 over 3

OpenStudy (anonymous):

\[\frac{ 32 + 4 * 4 -6 }{ 3 }\] like that?

OpenStudy (anonymous):

It's supposed to be 3 squared +(6-2)*4-6 over 3

OpenStudy (anonymous):

\[\frac{ 3^2 + 4 * 4 -6 }{ 3 }\] look right now?

OpenStudy (anonymous):

3squared + (6 − 2) • 4 − 6 over 3 That's how it is supposed to look

OpenStudy (anonymous):

6-2 = 4...so 3^2 + 4 * 4 - 6 divided by 3?